The statistical society claims the percent of people proficient at statistics is 15%. We sample 450 people and find 54 proficient at statistics. Find the z-score used in hypothesis testing. Standardized test statictics.
Answer by rothauserc(4718)

since we are looking for a z-score, we are dealing with a normal distribution
:
the mean of the population is the mean of the sample, 0.15
:
54/450 = 0.12
:
sample standard error = square root(0.15 * 0.85 / 450) = 0.0168
:
z-score = (0.12 - 0.15) / 0.0168 = −1.7857
: